Could you please help me out on this one?
Ron ateends a party. He wants to limit his food intake to 103g protein, 93g fat, and 135g carbohydrate. The mushroom caps at the party have 3g protein, 5g fat, and 9g carbs, the meatballs have 14g protein, 7g fat, and 15g carbs, and the deviled eggs have 13g protein, 15g fat, and 6g carbs.How many of each snack can he eat to obtain his goal?

Answer by stanbon(75887)

Ron ateends a party.
He wants to limit his food intake to 103g protein, 93g fat, and 135g carbohydrate.
The mushroom caps at the party have 3g protein, 5g fat, and 9g carbs,
the meatballs have 14g protein, 7g fat, and 15g carbs, and
the deviled eggs have 13g protein, 15g fat, and 6g carbs.
How many of each snack can he eat to obtain his goal?
----------------------
Let c be # of mushroom; m be # of meatballs; e be # of eggs
--------------------------------------
3c+14m+13e = 103
5c+ 7m+15e = 93
9c+15m+ 6e =135
------------------
Using the TI-83 Matrix function I get:
c=7
m=4
e=2
=============
Cheers,
Stan H.